The left bank is offering CD’s at 5%. The right bank is offering a money market account at 4%. If I have a total of $10,000 to invest in these accounts and want to earn $436 in interest at the end of one year, how much should I put into each account?
Answer by dkppathak(439)

The left bank is offering CD’s at 5%. The right bank is offering a money market account at 4%. If I have a total of $10,000 to invest in these accounts and want to earn $436 in interest at the end of one year, how much should I put into each account?
solution
Let left bank investment is $x
and right bank investment $10000-x
interest both account $436=x*0.05+0.04*(10000-x)
436=0.05x-0.04x +400
436-400=0.01x
36=0.01x
x=36/0.01=$3600
Answer
Left bank investment $3600 and right bank investment is $6400